The sum of two consecutive odd numbers is divisible by 4. Verify this statement with the help of some examples.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(i)5 + 3 = 8 which is divisible by 4
(ii) 7 + 9 = 16 which is divisible by 4
(iii) 13 + 15 = 28 which is divisible by 4

Thanks